OpenStudy (luigi0210):

Derivatives:

OpenStudy (luigi0210):

Find the 45th derivative of: \[\LARGE y=cos(2x)\]

zepdrix (zepdrix):

Ooo interesting problem :o

OpenStudy (nincompoop):

laughing out loud here's the trick, at the 4th derivative it goes back to original function

OpenStudy (nincompoop):

if you don't believe me, try it

zepdrix (zepdrix):

Well the sign/cosine go back to the start. We keep getting factors of 2 though. So it won't be exactly the same :)

OpenStudy (luigi0210):

Yup, it's that 2 that's the problem~

zepdrix (zepdrix):

So the 0th derivative has a 2^0 in front, yes?

zepdrix (zepdrix):

How bout the first derivative? 2^1?

zepdrix (zepdrix):

What do you think for the 45th? ;D

OpenStudy (luigi0210):

Don't you have to bring the 2 out in front tho? o_O

zepdrix (zepdrix):

\[\Large\bf\sf y0=\cos(2x)\]\[\Large\bf\sf y1=2(-\sin(2x))\]\[\Large\bf\sf y2=2^2(-\cos(2x))\]

zepdrix (zepdrix):

What do you mean igi? :o

OpenStudy (nincompoop):

you would put something in the front blah blah blah sin(2x)

OpenStudy (nincompoop):

-2^45 sin(2x)

OpenStudy (luigi0210):

Whoops, I mis-interpreted you guys. What I was trying to say is what you and nin just said

zepdrix (zepdrix):

lameeee +_+ you gave the answer.

OpenStudy (nincompoop):

laughing out loud

OpenStudy (luigi0210):

At least I know how he got it tho :3

zepdrix (zepdrix):

ok good c:

OpenStudy (nincompoop):

I usually don't do this, but I don't want him to actually put 2x2x2x2.....

zepdrix (zepdrix):

Do you though? c: Do you understand what he was saying with the multiples of 4 thing?

OpenStudy (luigi0210):

Yea, nin saved me time xD

OpenStudy (nincompoop):

someone is late in the game

OpenStudy (luigi0210):

Thanks guys :)